OpenStudy (anonymous):

t:P2-->P2 by x^2->x^2+1, x->x^2-1, and 1->3 how would you find T=RepB,B(t) with B=

OpenStudy (anonymous):

You have the basis {1, x, x^2}, right? T(1) =3 T(x) = -1 +x^2 T(x^2) = 1+x^2 so that T=\(\left[\begin{matrix}3&-1&1\\0&0&0\\0&1&1\end{matrix}\right]\)

OpenStudy (anonymous):

got what I mean?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

good

OpenStudy (anonymous):

that would be the final answer of the matrix correct?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so to find S where D=<1,1+x,1+x+x^2> would T=

OpenStudy (anonymous):

T= 1 1 1 0 1 1 0 0 1

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

from there how do you fin matrix p=P^-1SP

OpenStudy (anonymous):

?